Suppose that when Mary’s income increases from RM56,000 to RM60,000 per year, her

purchase of good B decreased from 10 units to 5 units.

i. Calculate the coefficient of income elasticity of demand.

ii. Interpret the value of the coefficient above.


1
Expert's answer
2022-02-17T11:23:15-0500

i) percentage change in income ="\frac{60000-56000}{56000}\times 100\%=7.143\%"

Percentage change in quantity demanded "=\frac{5-10}{10}\times 100\%=-50\%"

Income elasticity of demand "=" percentage change in quantity demanded"\/" percentage change in income"=\frac{7.143}{-50}=-0.14286"


ii) The value of the coefficient is negative showing that good B is an inferior good.